40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 9205     Entry time: Sun Oct 6 17:05:49 2013
Author: rana 
Type: Summary 
Category: IOO 
Subject: MC ASC problems 

MC unlocked over the weekend and also got severely mis-aligned. It all started around midnight on Saturday.

At first I thought that this was due to the MCS CPU meter being railed at 60 us, so I deleted a bunch of filters in MC1,2,3 that are unused and left over from Den's quantization noise investigations. This reduced the CPU load somewhat, but didn't make any real improvements. Turning on the ASC filter banks in the MC SUS still mis-aligned the MC.

With the MC WFS and MC ASS turned off, there is still some digital junk coming in and misaligning things. Plot attached.

Similar stuff coming in on ITMX, but not ITMY.

Tried restarting various FEs, but there was no effect. Also tried rebooting c1lsc, c1ioo, & c1sus. Finally did 'shutdown -r now' on all 5 computers on the CDS overview screen and simultaneously (almost) pressed the reset button on the RFM switch above the old c1pem crate. Everything came back OK except for c1oaf (I had to manually button his BURT button) and now the ASC inputs on all the SUS are zero when they should be and MC is well locked and aligned.

Rob and I used to do this trick when he thought that a cosmic ray had corrupted a bit in the RFM network.

Attachment 1: mcasc.pdf  10 kB  | Hide | Hide all
mcasc.pdf
ELOG V3.1.3-